OpenStudy (anonymous):

20 IS 4% OF WHAT?? PLZ ANSWER

OpenStudy (paki):

20=4/100*X.... solve for "x"....

OpenStudy (texaschic101):

20 is 4% of what... " is " means equals " of " means multiply 4% turned to decimal is .04 " what " = x 20 = .04x -- now divide both sides by .04 and you will have your answer

OpenStudy (anonymous):

I JUST NEED THE ANSWER ON A TIMED TEST

OpenStudy (anonymous):

HELP PLZZZ

OpenStudy (texaschic101):

500

OpenStudy (anonymous):

What percent discount did Josh get when he saved two dollars on a shirt that was originally sixteen dollars? eight percent twelve point five percent twenty−five percent thirty−two percent

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

PLEASE HELP ME WITH THIS

OpenStudy (texaschic101):

$2 is what % of 16 percent formula : is/of = %/100 2/16 = x/100 cross multiply 16x = 200 x = 12.5%